# Revisiting Locally Differentially Private Protocols: Towards Better Trade-offs in Privacy, Utility, and Attack Resistance

### v2

@@ VI-D Pareto Frontier for ASR and MSE
ASR-MSE trade-offs are protocol-dependent: Our findings show that each protocol exhibits a distinct ASR-MSE profile. By jointly optimizing for both ASR and MSE, our adaptive protocols consistently push these curves toward more favorable regimes in the ASR-MSE Pareto frontier.
VI-E Impact of Weights in the Objective Function
Thus far, we have focused on analyzing the performance of adaptive protocols under fixed objective configurations for the weights $(w_{\text{ASR}}=w_{\text{MSE}}=0.5)$ in Equation (23). However, our proposed multi-objective framework introduces a new degree of freedom: practitioners can adjust the relative importance of ASR vs. variance (MSE) when optimizing LDP protocols. Figure 6 illustrates how varying these weight combinations $(w_{\text{ASR}},w_{\text{MSE}})$ influences the ASR, MSE, and optimal parameter choices for AUE ($p$), ALH ($g$), ASS ($\omega$) and ATHE ($\theta$) under a fixed setting ($k=100$, $\varepsilon=4$).
Figure 6: Optimal parameter choices for each adaptive protocol as a function of the weight combination $(w_{\text{ASR}},w_{\text{MSE}})$, evaluated at $k=100$ and $\varepsilon=4$. Each sub-figure compares the adaptive protocol’s chosen parameter (blue color curve) against the corresponding parameter choice in the original, non-adaptive protocol (red color dashed line).
From Figure 6, one can notice that as the weight on ASR $w_{\text{ASR}}$ increases, each adaptive protocol tends to choose parameter values that more aggressively reduce the ASR at the cost of increasing the MSE. Conversely, placing greater emphasis on MSE drives parameters toward configurations closer to or equal to those of the original protocols (i.e., SS, OUE, OLH and THE), aiming to preserve utility even if it elevates the ASR. These results confirm the benefits of our two-objective optimization framework: rather than a static parameter choice, practitioners can tune the protocol parameters in response to changing priorities, achieving a more flexible trade-off between privacy (ASR) and utility (MSE).
VI-F Adaptive and Optimized Parameters
In this section, we analyze the optimization of parameters in our adaptive protocols (AUE, ALH, ASS and ATHE) by examining the behavior of their objective functions (Equations (24)–(27)), which balance the ASR-MSE trade-off. Figure 7 illustrates the objective function as a function of key parameters: AUE ($p$), ALH ($g$), ASS ($\omega$) and ATHE ($\theta$), with a fixed $k=100$ and $\varepsilon=4$. The selected parameters for our adaptive protocols are compared against the fixed state-of-the-art parameters of OUE, OLH, SS and THE.
Figure 7: Objective function value as a function of key parameters for our adaptive protocols (AUE, ALH, ASS and ATHE) compared with their state-of-the-art counterparts (OUE, OLH, SS and THE). Vertical dashed lines (green color) indicate the optimal parameter values selected by our adaptive protocols, while vertical dash-dotted lines (red color) represent the fixed parameter values of the state-of-the-art protocols. Without loss of generality, we set $k=100$ and $\varepsilon=4$.
Figure 8: Comparison of empirical and analytical Pareto frontiers for ASR vs. Variance (MSE) across various LDP protocols (state-of-the-art and adaptive). Each subplot considers a range of privacy budgets $\varepsilon\in(2,10)$ and a fixed domain size $k=100$. Empirical results ($\circ$ markers) are averaged over $100$ independent runs with the Age attribute of the Adult dataset [46], while analytical results (red dashed lines) are computed using closed-form equations.
For AUE, we observe in the top-left plot that the objective function reaches its minimum at $p=0.818$, which is notably higher than the fixed $p=0.5$ used by OUE. This also means that parameter $q$ will increase to satisfy $\varepsilon$-LDP, i.e., increasing the probability of reporting random bits. For ALH, as shown in the top-right plot, the optimal hash domain size is reduced to $g=13$ in our adaptive protocol compared to $g=56$ in OLH. This reduction in $g$ lowers the ASR at the expense of slightly increased variance, aligning with our adaptive objective to achieve a better balance between privacy and utility. For ASS, as depicted in the bottom-left plot, the optimal subset size $\omega=7$ contrasts with the fixed $\omega=2$ used in SS. The larger $\omega$ effectively spreads the probability mass across a larger subset, reducing ASR while incurring a higher variance. For ATHE, the bottom-right plot shows that the adaptive threshold $\theta=0.783$ is slightly lower than the fixed threshold $\theta=0.816$ in THE. This subtle adjustment enables ATHE to reduce privacy attacks while maintaining competitive variance levels, showcasing the precision of our adaptive optimization.
Adaptive protocols optimize parameters for better trade-offs: Our findings demonstrate that the optimized parameters selected by adaptive protocols significantly differ from the fixed parameters of state-of-the-art protocols, resulting in improved robustness to privacy attacks with controlled increases in variance. This optimization highlights the effectiveness of our adaptive mechanisms in better navigating the privacy-utility trade-off space.
VI-G Empirical Pareto Frontier for ASR and MSE
To assess the accuracy of our closed-form ASR and MSE equations within the ASR-MSE two-objective framework (Section V-B), we conduct empirical experiments using the Adult dataset from the UCI machine learning repository [46]. We select the Age attribute with domain size $k=100$ (i.e., $\text{Age}\in[0,99]$) and $n=48842$ users, reporting results averaged over $100$ independent runs.
Figure 8 presents the comparison between empirical ($\circ$ markers) and analytical (red dashed lines) Pareto frontiers for ASR versus MSE across various LDP protocols, including both state-of-the-art and our adaptive variants. Each subplot represents a specific protocol (i.e., GRR, SUE, BLH, SHE, SS, OUE, OLH, THE, ASS, AUE, ALH and ATHE) with the privacy budget varying as $\varepsilon\in(2,10)$ with the domain size fixed at $k=100$. One can notice that across all protocols, the empirical results closely align with the analytical calculation, showcasing the robustness of the closed-form equations for ASR and MSE. This alignment highlights the reliability of our theoretical framework and previous analytical experiments across different privacy budgets and data distributions. To further support this claim, similar to Figure 1, Appendix -D provides additional experiments using a synthetic dataset generated from a Dirichlet distribution ($\mathbf{1}$), exploring variations in the number of users ($n\in\{5000,50000,500000\}$).
Analytical vs. empirical validation: Our findings demonstrate that the analytical results closely align with empirical ones, reaffirming the reliability of the closed-form equations across various privacy budgets and scenarios.
@@ VII Conclusion and Perspectives
In this work, we introduced a general multi-objective optimization framework for refining LDP frequency estimation protocols under adversarial conditions. Unlike prior approaches that minimize estimation error (MSE) in isolation, our framework allows for principled, flexible trade-offs between utility and privacy, selecting protocol parameters through classical multi-objective optimization strategies while strictly preserving the original $\varepsilon$-LDP guarantees. As an instantiation of this framework, we focused on a two-objective formulation that jointly minimizes the Attacker Success Rate (ASR) under data reconstruction attacks [28, 26, 29] and Mean Squared Error (MSE), demonstrating that existing protocols (i.e., SS, OUE, OLH and THE) can be substantially improved through principled parameter tuning. Our adaptive protocols, namely, ASS, AUE, ALH and ATHE, can significantly reduce privacy leakage while maintaining high accuracy across a wide range of scenarios, making our methods directly applicable in real-world systems.
To our knowledge, this is the first work to jointly optimize analytical ASR and MSE via flexible, selector-driven parameter tuning, all while preserving $\varepsilon$-LDP and extending analytical formulations to new protocols. This contribution bridges a gap between theoretical guarantees and practical deployment risks, offering a robust methodology to assess and mitigate vulnerabilities in LDP mechanisms. Looking ahead, our framework is extensible to broader privacy, utility, and security objectives, including robustness to poisoning [34, 36], resistance to inference [27, 9, 33] or re-identification attacks [25, 26], and even communication efficiency. Finally, we envision this work as a foundation for a modular LDP optimization suite, paving the way for adaptive, attack-aware LDP deployments at scale.
Acknowledgment
This work has been partially supported by the French National Research Agency (ANR), under contract “ANR-24-CE23-6239” JCJC project AI-PULSE and by the “ANR 22-PECY-0002” IPOP (Interdisciplinary Project on Privacy) project of the Cybersecurity PEPR. Sébastien Gambs is supported by the Canada Research Chair program as well as a Discovery Grant from NSERC.

### v3

@@ VI-D Pareto Frontier for ASR and MSE
Key insight: Each protocol exhibits a distinct ASR-MSE profile. By jointly optimizing for both ASR and MSE, our adaptive protocols consistently push these curves toward more favorable regimes in the ASR-MSE Pareto frontier.
VI-E Complementary Adversarial Perspective: Poisoning Attacks
While ASR captures vulnerability to data reconstruction attacks, LDP protocols may also be susceptible to data poisoning attacks [34], in which an adversary injects malicious reports to bias the aggregated frequency estimates. To capture this integrity-oriented threat (see Eq. (26)), we adopt the maximal gain attack (MGA) proposed by Cao et al. [34], which is shown to be the strongest poisoning strategy in terms of maximizing the expected increase in the estimated total frequency of a target item set $T$. For UE and LH protocols, the MGA admits a unified closed-form expression, referred to as the attacker’s Expected Poisoning Gain (EPG):
$\mathrm{EPG}=\beta\!\left(\frac{r(1-q)}{p-q}-f_{T}\right),$ (32)
in which $\beta$ denotes the fraction of fake users, $r=|T|$ is the number of targeted items and $f_{T}$ is the total true frequency mass of these items among genuine users. Following the experimental setup of [34], we fix the poisoning parameters to $\beta=0.05$, $r=1$, and $f_{T}=0.01$, ensuring comparability and avoiding additional attacker-side tuning degrees of freedom.
Joint View. Taken together, ASR, EPG and MSE define a three-dimensional trade-off between privacy leakage, estimation utility and integrity robustness. Figure 6 provides a geometric illustration of this joint space for state-of-the-art UE- and LH-based protocols and our adaptive counterparts, where each point corresponds to a distinct $\varepsilon\in(2,10)$ and a fixed $k=100$. A key implication is that decreasing $\varepsilon$ simultaneously increases the estimation error (MSE) and poisoning vulnerability (EPG), while reducing the reconstruction risk (ASR). Rather than prescribing a single operating point, this three-objective view supports multi-objective optimization or projections onto lower-dimensional trade-offs [47] (e.g., ASR vs. MSE). In our experiments, adaptive protocols are tuned exclusively with respect to ASR and MSE (Section V-C), while EPG is evaluated post hoc to assess the resulting integrity implications. More generally, alternative adaptive designs could directly incorporate poisoning robustness into the optimization objective, e.g., through weighted combinations of ASR, MSE and EPG as Eq. (27), which we leave as future work.
(a) UE-based protocols: SUE, OUE, and AUE.
(b) LH-based protocols: BLH, OLH, and ALH.
Figure 6: Joint ASR-MSE-EPG trade-off for UE- and LH-based LDP frequency estimation protocols at domain size $k=100$. Each point corresponds to a distinct privacy budget $\varepsilon\in(2,10)$.
Within this three-dimensional space, our adaptive protocols (AUE and ALH) reshape the attainable trade-offs at fixed $\varepsilon$. For UE-based protocols, AUE systematically bridges the gap between the utility-oriented OUE and the more conservative SUE, achieving comparable reconstruction rates to SUE at substantially lower estimation error. For LH-based protocols, ALH reduces ASR relative to OLH and, as a consequence of its parameter adaptation, also yields lower EPG, while achieving robustness levels comparable to BLH with substantially improved utility. Overall, these results demonstrate that attack-aware parameter adaptation reveals operating points that cannot be achieved by tuning the privacy budget alone.
@@ VII Conclusion and Perspectives
This work presented a comprehensive experimental and analytical benchmark of LDP frequency estimation protocols, emphasizing the joint evaluation of privacy leakage, utility, and integrity under adversarial settings. By positioning existing protocols within a unified privacy-utility-attackability trade-off space, we provided a reproducible methodology to assess and compare their resilience against both data reconstruction and poisoning attacks. Our study revealed clear and protocol-dependent Pareto frontiers between ASR under data reconstruction attacks [30, 28, 31] and MSE, as well as their interaction with integrity-oriented attackability, exposing vulnerabilities not captured by $\varepsilon$ alone.
Beyond analysis, we showed that systematic benchmarking can guide the refinement of existing mechanisms. By leveraging multi-objective optimization, we derived adaptive parameterizations (ASS, AUE, ALH, ATHE) that achieve more favorable privacy-utility-attackability trade-offs than their original counterparts (SS, OUE, OLH, THE) while strictly preserving $\varepsilon$-LDP guarantees. These adaptive variants demonstrate that informed tuning can substantially enhance robustness against adversarial threats without prohibitive utility loss.
Overall, our work establishes a principled approach to evaluating LDP mechanisms through standardized, attack-aware benchmarks that jointly capture privacy, utility and integrity considerations. We envision this benchmark as a foundation for a broader suite of experimental tools for LDP, enabling future studies to extend our evaluation to additional adversarial models (e.g., attribute inference [29, 10, 35], re-identification [27, 28]) and complementary system-level metrics such as communication and computation cost.
Acknowledgments
The authors thank the anonymous reviewers for their insightful suggestions and Haoying Zhang for her helpful comments on an earlier draft. Héber H. Arcolezi has been partially supported by the French National Research Agency (ANR), under contracts: “ANR-24-CE23-6239” and “ANR-23-IACL-0006”. Sébastien Gambs is supported by the Canada Research Chair program as well as a Discovery Grant from NSERC.
AI-Generated Content Acknowledgement
The authors acknowledge the use of ChatGPT (OpenAI, GPT-5 model) to assist with language-related improvements, including grammar correction, spelling, formatting consistency and refinement of phrasing for clarity and readability. No part of the paper’s scientific content, analysis, experimental design, results or conclusions was generated or influenced by AI tools. All conceptual, technical and experimental contributions originate solely from the authors.